# What’s the Opposite of a Face? Finding Shared Decodable Concepts and their Negations in the Brain

### v1

@@ 2.4 Decoding Methodology
$\text{Contrast}(\bm{A},\bm{B})=-\frac{1}{M}\sum_{i=1}^{M}log\bigg{(}\frac{exp(d(\bm{a}_{i},\bm{b}_{i})/\tau)}{\sum_{j=1}^{M}exp(d(\bm{a}_{i},\bm{b}_{j})/\tau)}\bigg{)}$ (1)
$\mathcal{L}_{\text{contrastive}}(\bm{A},\bm{B})=\frac{1}{2}[\text{Contrast}(\bm{A},\bm{B})+\text{Contrast}(\bm{B},\bm{A})]$ (2)
In this definition $\bm{A}=[\bm{a}_{1},\ldots,\bm{a}_{n}]$ and $\bm{B}=[\bm{b}_{1},\ldots,\bm{b}_{n}]$ are embeddings for two modalities representing the same data points, $d$ is cosine distance, and $\tau$ is a temperature hyper-parameter. The loss is minimized when the distance between matching embeddings is small, and the distance between mismatched embeddings is large. In the original CLIP setting, $\bm{A}$ and $\bm{B}$ represent embeddings for images and corresponding text captions. In our implementation, we apply the contrastive loss to image embeddings computed by a pretrained frozen CLIP model and CLIP embeddings predicted from fMRI, i.e. we optimize $\min_{\bm{\theta}}\>\mathcal{L}(\bm{\hat{Y}},\bm{Y}_{\text{CLIP}})=\mathcal{L}(g(\bm{X};\bm{\theta}),\bm{Y})$. An illustration of the decoding procedure is given in Figure 11 (additional details on the decoding procedure are given in Section 6.1).
@@ 3 Finding Shared Brain-Decodable Concepts and their Representative Images
In this section, we describe our methods for finding and interpreting areas of the brain responsive to semantically similar images across participants. We approach this problem by analyzing the weight matrices $\bm{W}^{(k)}$ from the optimized linear decoders $g^{(k)}(\bm{x})=\bm{W}^{(k)}\bm{x}=\bm{y}$ for all subjects $k\in\{1\ldots 8\}$. First, we observe that the decoder’s linear transformation $\bm{W}^{(k)}\bm{x}=\sum_{i=1}^{v}\bm{w}_{i}^{(k)}\cdot{x_{i}}$ is simply a summation of parameter vectors $\bm{w}_{i}^{(k)}\in\mathbb{R}^{512}$ that are scaled by brain response values $x_{i}\in\mathbb{R}$. The key to our analysis is to view the parameter vector $\bm{w}_{i}^{(k)}$ as a CLIP vector that represents a brain-decodable concept for voxel $i$. The allows us to use cosine distance $d(\bm{w}_{i}^{(k)},\bm{w}_{j}^{(r)})$ between voxels $i,j$ from participants $k,r$ as a measure of similarity between the decodable concepts of brain voxels. A schematic of the algorithm we apply is given in Figure 1.
@@ 3.1 Cross Participant Clustering
To discover shared brain-decodable concepts we apply a novel clustering method to the per-voxel model parameter vectors $\bm{w}_{i}^{(k)}$ across all participants. To reduce the impact of random initialization and combat the noise inherent in fMRI, we re-train the decoder model 50 times for every participant and compute the per-participant average of the resulting linear decoder weight matrices.
@@
We base our clustering method on the density-based spatial clustering of applications with noise (DBSCAN) (Ester et al., 1996) algorithm, which is parameterized by a neighborhood size $\varepsilon$ and a point threshold minPts. The algorithm can be summarized using the following steps: (1) Points that have at least minPts points in their $\varepsilon$-neighborhood are marked as core points; (2) construct a graph $\mathcal{G}$ where the vertices are core points, and there are edges between core points that have a distance less than $\varepsilon$; (3) clusters are formed by finding the connected components of $\mathcal{G}$ and (4) the remaining non-core points are added to clusters if they are within the $\varepsilon$-neighborhood of a core point, otherwise, they are marked as outliers not belonging to any cluster. Figure 2 outlines the application of our modified DBSCAN algorithm to modeling fMRI data.
@@
In our modified cross-participant DBSCAN, we redefine the core point threshold minPts as minNeighbors. Then, in step (1) a point $\bm{w}_{i}^{(k)}$ is a core point if there are points from at least minNeighbors other participants within its $\varepsilon$ neighborhood. Since there are 8 participants in the study, the maximum value of minNeighbors is 7. Steps (2) to (4) proceed as above.
@@
Our second modification is to apply a within-participant expansion of clusters. We define a new hyperparameter $\varepsilon_{\text{expansion}}$ and apply a final expansion step: (5) All points inside the $\varepsilon_{\text{expansion}}$ neighborhood of a point in a cluster become members of that cluster with the constraint that they must be from the same participant. Unlike steps 1-3, it is possible for a point to be assigned to multiple clusters in step (5). We present results using a fixed value for $\texttt{minNeighbors}=3$, with a range of values for $\varepsilon\in\{0.5,0.55,0.6,0.65\}$ to explore clusters at different density scales. We set the expansion neighborhood size $\varepsilon_{\text{expansion}}=\text{min}(\varepsilon+0.05,0.65)$ so that the cluster boundaries can grow slightly larger. The number of cross-participant clusters found was 9, 14, 12, 11 for $\varepsilon$ values of 0.5, 0.55, 0.6, 0.65 respectively.
@@ 3.2 Selecting Representative Images
To interpret the semantic meaning of a vector $\bm{w}$ in CLIP space, we can consider the nearby images. To select nearby images we find the cosine distance $D(\bm{w},\bm{Y})=\{d(\bm{w},\bm{y})|\bm{y}\in\bm{Y}\}$ to the set of CLIP image embeddings $\bm{Y}$. The image embeddings that have the smallest distance to $\bm{w}$ are chosen as the positive representative data points, and we display their corresponding images. Similarly, we use $D(-\bm{w},\bm{Y})$ to find a set of images that are associated with the negated vector, which often represents a concept that is opposite to $\bm{w}$. We denote these as the negative representative images.
In order to find the representative images for an SDC cluster, we first computed the cluster centroid by taking the mean of all parameter vectors within the cluster $\bm{\bar{w}}=\frac{1}{|I|}\sum_{(i,k)\in I}\bm{w}_{i}^{(k)}$ where $I$ is a set of tuples identifying the voxel and subject indices for an SDC cluster. DBSCAN has the ability to create non-globular clusters for which the centroid might not be a good representative. In our experience the centroid is within $\varepsilon$ of a core point within the cluster, implying that it is likely fairly representative of the cluster.
Instead of using the original CLIP embeddings for each image, we instead used the subject-specific brain-decoded embeddings $\bm{\hat{Y}}_{\text{Test}}^{\text{k}}$. This allows us to focus specifically on the information that can be decoded from fMRI, which could be a subset of the information represented by CLIP space. For images that were viewed more than once we average the embeddings within and across participants. These averaged brain-decoded embeddings from held-out test data are denoted by $\bm{\hat{Y}}_{\text{Test}}^{\text{AVG}}$. Finally, we take the cosine distances between the SDC cluster centroid $D(\bm{\bar{w}},\bm{\hat{Y}}_{\text{Test}}^{\text{AVG}})$ and the negated centroid $D(-\bm{\bar{w}},\bm{\hat{Y}}_{\text{Test}}^{\text{AVG}})$. We select the images with the smallest distances to $\bm{\bar{w}}$ and $-\bm{\bar{w}}$ as the positive and negative representative images respectively.
@@ 4 Results
Our clustering method not only identifies previously discovered functional areas but additionally identifies new ones, which we explore in this section. We note that there is no constraint that voxels within a shared visual semantic cluster be spatially adjacent, yet we consistently find contiguous patches of voxels during visualization of our clusters on the flattened cortex brain maps both within and across participants. We can localize these patches of voxels to specific regions of interest (ROIs) both functionally and anatomically (See Section 6.3 for further details).
@@ 4.1 Faces
One of the first reported functionally localized areas for higher-order vision was the fusiform face area (FFA)  Kanwisher et al. (1997). Our method identifies a face-related concept (Figure 4.1 $\varepsilon=0.55$, cluster 7) that is localized to FFA and includes voxels from all 8 participants. We also see some voxels in the extrastriate body area (EBA), likely because many face images include part or all of a person’s body. We note that the positive representative images are not exclusively human and include a range of animal faces. The images often depict people eating and handling food, or holding other objects such as toothbrushes or cell phones.
@@ 4.3 Bodies
Our method identifies five notable body-related areas at $\varepsilon=0.55$ in and around EBA. The positive representative images for cluster 2 (Figure 5) show people and animals outside with an emphasis on legs and active movement. The negative images typically depict people indoors sitting with their legs obscured by tables. Cluster 11 (Figure 6) has a similar emphasis on hands instead of legs. People are displayed in a variety of contexts with their hands clearly visible, while the negative images are exclusively non-primate animals without hands. This cluster also shows indoor/outdoor contrast in the representative image groups and so there is some PPA activation. In Section 6.7 we outline further body-related shared voxel clusters associated with motion and a distinction between single individuals and groups of people in a sub-region of EBA.
@@ 5 Conclusion
We proposed a novel mechanism to identify shared decodable concept clusters in a large multi-participant fMRI dataset of natural image viewing (Allen et al., 2022). To achieve this, we applied a contrastive learning approach using multimodal embeddings from CLIP (Radford et al., 2021) and a novel adaptation of the DBSCAN clustering algorithm (Ester et al., 1996). Our method dentifies previously reported functional areas that code for various visuo-semantic categories, as well as identifying novel shared categories. Furthermore, our method exposes interpretability by identifying the most / least associated images with CLIP-based cluster centroids. In the case of faces, we find a set of positive / negative images where the negative images depict scenarios where a face might be expected, but is not visible. This suggests that the brain responses for this SDC cluster are organized in a way that specifically codes for the presence of faces in expected visual depictions. Food is a concept that is highly confounded with other visual criteria such as shape and color. By providing the most negatively associated images for a SDC cluster, we enabled easy interpretation of semantic coding, e.g. in the case where the negated images to this SDC cluster are all grayscale images (see Section 4.2). We further identify SDC clusters specific to bodies in motion, in groups, as well as hands and legs. We identify SDC clusters related to repeated elements (numerosity), sensitivity to indoor / outdoor scenes, cardinal orientation of objects in images (horizontal / vertical) and gradations of shape information (soft vs hard) and visual perspective (close vs far). Our method showcases how various techniques from machine learning can be employed to better characterize visual representations of semantic selectivity in the human brain.
5.1 Limitations
If concepts are not encoded by the BOLD signal in fMRI, we will not be able to recover them. On the other hand, the methods we develop here could be applied to other modalities that more directly measure neural activity (electrocorticography). Consequently, this first limitation is one that could be overcome in future studies. Secondly, we are also only able to uncover concepts that appear in the CLIP space spanned by the images in the stimulus set. This means that, while our method is more flexible than the hypothesis-driven ones used by previous studies (Kanwisher et al., 1997), some important concept representations could still be missed by our method. Future work could address this limitation by using even larger stimulus sets. Finally, we report on a few probable localizations of concepts. These provide clear hypotheses for future experiments that should be targeted at attempting to confirm these findings. This emphasizes that our new data-driven method is not a replacement for the hypothesis-driven approach to identifying concept-encoding regions in the human brain: rather, it serves as a data-driven hypothesis generator.
Broader Impacts
Methods that decode brain activity could have substantial impacts on society. These methods could assist in diagnosing psychiatric disorders, or in helping individuals with locked-in syndrome or related disorders to better communicate. However, the same methods pose privacy concerns. As with all emerging technologies, responsible deployment is needed in order for society as a whole to obtain maximum benefit while mitigating risk.

### v2

# Finding Shared Decodable Concepts and their Negations in the Brain

@@ 2.4 Decoding Methodology
$\text{Contrast}(\bm{A},\bm{B})=-\frac{1}{M}\sum_{i=1}^{M}log\bigg{(}\frac{exp(\bm{a}_{i}\cdot\bm{b}_{i}/\tau)}{\sum_{j=1}^{M}exp(\bm{a}_{i}\cdot\bm{b}_{j}/\tau)}\bigg{)}$ (1)
$\mathcal{L}_{\text{InfoNCE}}(\bm{A},\bm{B})=\frac{1}{2}[\text{Contrast}(\bm{A},\bm{B})+\text{Contrast}(\bm{B},\bm{A})]$ (2)
In this definition $\bm{A}=[\bm{a}_{1},\ldots,\bm{a}_{n}]$ and $\bm{B}=[\bm{b}_{1},\ldots,\bm{b}_{n}]$ are embeddings for two modalities representing the same data points, the $\cdot$ operator represents cosine similarity, and $\tau$ is a temperature hyper-parameter. The loss is minimized when the distance between matching embeddings is small, and the distance between mismatched embeddings is large. In the original CLIP setting, $\bm{A}$ and $\bm{B}$ represent embeddings for images and corresponding text captions. In our implementation, we apply the contrastive loss to image embeddings computed by a pretrained frozen CLIP model and CLIP embeddings predicted from fMRI, i.e. we optimize $\min_{\bm{\theta}}\>\mathcal{L}(\bm{\hat{Y}},\bm{Y}_{\text{CLIP}})=\mathcal{L}(g(\bm{X};\bm{\theta}),\bm{Y})$. An illustration of the decoding procedure is given in Figure 11 (additional details on the decoding procedure are given in Section 6.1).
@@ 3 Finding Shared Brain-Decodable Concepts and their Representative Images
In this section, we describe our methods for finding and interpreting areas of the brain that respond to semantically similar images, and that are consistent across participants. We approach this problem by analyzing the weight matrices $\bm{W}^{(k)}$ from the optimized linear decoders described in section 3.4 $g^{(k)}(\bm{x}^{(k)})=\bm{W}^{(k)}\bm{x}^{(k)}=\bm{y}$ for all subjects $k\in\{1\ldots 8\}$. First, we observe that the decoder’s linear transformation $\bm{W}^{(k)}\bm{x}^{(k)}=\sum_{i=1}^{v}\bm{w}_{i}^{(k)}\cdot{x_{i}}^{(k)}$ is simply a summation of parameter vectors $\bm{w}_{i}^{(k)}\in\mathbb{R}^{512}$ that are scaled by brain response values $x_{i}^{(k)}\in\mathbb{R}$. This means that when $x_{i}^{(k)}$ is above or below baseline activation, the CLIP dimension represented by $\bm{w}_{i}^{(k)}$ is increased or decreased in the brain-decoded embedding. This motivates the key to our analysis: we view the parameter vector $\bm{w}_{i}^{(k)}$ as a CLIP vector that represents a brain-decodable concept for voxel $i$. This allows us to use cosine distance $d(\bm{w}_{i}^{(k)},\bm{w}_{j}^{(r)})$ between the weight vectors for voxels $i,j$ from participants $k,r$ as a measure of similarity between the decodable concepts of brain voxels across participants. Our objective is to apply a clustering algorithm using this metric in order to find areas in the brains of multiple participants that have a high conceptual similarity. We refer to these underlying concepts as shared decodable concepts (SDCs). We interpret the SDCs by retrieving a set of representative images associated with the brain-decoded CLIP vectors that are closest to the centroid of each SDC cluster. A schematic of the algorithm we apply is given in Figure 1.
@@ 3.1 Cross Participant Clustering
To discover shared brain-decodable concepts we apply a novel clustering method to the per-voxel model parameter vectors $\bm{w}_{i}^{(k)}$ across all participants. We base our clustering method on the density-based spatial clustering of applications with noise (DBSCAN) (Ester et al., 1996) algorithm, which is parameterized by a neighborhood size $\varepsilon$ and a point threshold minPts. The algorithm can be summarized using the following steps: (1) Points that have at least minPts points in their $\varepsilon$-neighborhood are marked as core points; (2) construct a graph $\mathcal{G}$ where the vertices are core points, and there are edges between core points that have a distance less than $\varepsilon$; (3) clusters are formed by finding the connected components of $\mathcal{G}$ and (4) the remaining non-core points are added to clusters if they are within the $\varepsilon$-neighborhood of a core point, otherwise, they are marked as outliers not belonging to any cluster.
@@
In our modified cross-participant DBSCAN, we redefine the core point threshold minPts as minNeighbors. Then, in step (1) a point $\bm{w}_{i}^{(k)}$ is a core point if there are points from at least minNeighbors other participants within its $\varepsilon$ neighborhood. With this modification, a core point now identifies voxels that represent a brain-decodable concept that is shared with at least minNeighbors participants. Since there are 8 participants in the study, the maximum value of minNeighbors is 7. Steps (2) to (4) proceed as above.
Our second modification is to apply a within-participant expansion of clusters. We define a new hyperparameter $\varepsilon_{\text{expansion}}$ and apply a final expansion step: (5) All points inside the $\varepsilon_{\text{expansion}}$ neighborhood of a point in a cluster become members of that cluster with the constraint that they must be from the same participant. The addition of this step was motivated by our early experiments, where we noticed that there were sometimes only one or two voxels belonging to certain participants for each cluster. This allows the cluster boundaries to grow slightly larger for each participant, and we found that this did not significantly change the cluster location or semantic interpretations. Figure 2 outlines the application of our modified DBSCAN algorithm to modeling fMRI data.
@@
We present results using a fixed value for $\texttt{minNeighbors}=3$. This means that each cluster must include at least 4 out of 8 participants in the study. This allows for the discovery of SDCs that may show up less reliably in some participants. Furthermore, we use a range of values for $\varepsilon\in\{0.5,0.55,0.6,0.65\}$ to explore clusters at different density scales. We set the expansion neighborhood size $\varepsilon_{\text{expansion}}=\text{min}(\varepsilon+0.05,0.65)$ so that the cluster boundaries can grow slightly larger than the baseline neighborhood size $\varepsilon$, but not exceeding 0.65 as we noticed neighborhoods become over-connected as $\varepsilon$ approaches 0.7. The number of cross-participant clusters found was 9, 14, 12, 11 for $\varepsilon$ values of 0.5, 0.55, 0.6, 0.65 respectively.
To reduce the impact of random initialization and combat the noise inherent in fMRI, we re-trained the decoder model 50 times for every participant and compute the average of the resulting parameter vectors. In other words, each averaged parameter vector $\bm{w}_{i}^{(k)}$ used in our analysis represents the average of 50 concepts that a voxel could represent depending on the random initialization of the decoder. We found that this significantly improved the quality and consistency of the clusters revealed by our modified DBSCAN algorithm.
@@ 3.2 Selecting Representative Images
To interpret the semantic meaning of an SDC cluster, we first computed the cluster centroid by taking the mean of all parameter vectors within the cluster $\bm{\bar{w}}=\frac{1}{|I|}\sum_{(i,k)\in I}\bm{w}_{i}^{(k)}$ where $I$ is a set of tuples identifying the voxel and subject indices for an SDC cluster. When the cluster voxels are above or below baseline activation, the CLIP dimension represented by $\bm{\bar{w}}$ is increased or decreased in the brain-decoded embedding, respectively. We note that DBSCAN has the ability to create non-globular clusters for which the centroid might not be a good representative. In our experience the centroid is within $\varepsilon$ of a core point within the cluster, implying that it is likely fairly representative of the cluster.
In order to find the representative images for an SDC cluster, we can consider the nearby image-embeddings to the cluster centroid $\bm{\bar{w}}$. Instead of using the original CLIP embeddings for each image, we instead used the subject-specific brain-decoded embeddings $\bm{\hat{Y}}_{\text{Test}}^{(k)}$. This allows us to focus specifically on the information that can be decoded from fMRI, which could be a subset of the information represented by CLIP space. For images that were viewed more than once we average the embeddings within and across participants. These averaged brain-decoded embeddings from held-out test data are denoted by $\bm{\hat{Y}}_{\text{Test}}^{\text{AVG}}$. To select nearby images we define $D(\bm{w},\bm{Y})=\{d(\bm{w},\bm{y})|\bm{y}\in\bm{Y}\}$ which is the set of distances between a CLIP vector $\bm{w}$ to a set of CLIP embeddings $\bm{Y}$. We take the cosine distances between the SDC cluster centroid $D(\bm{\bar{w}},\bm{\hat{Y}}_{\text{Test}}^{\text{AVG}})$ and the negated centroid $D(-\bm{\bar{w}},\bm{\hat{Y}}_{\text{Test}}^{\text{AVG}})$ and retrieve the images corresponding to the smallest values in these sets as the positive and negative representative images respectively.
@@ 4 Results
The clustering method described in chapter 3 not only identifies previously discovered functional areas but additionally identifies new ones, both of which we explore in this section. We emphasize that there is no constraint that voxels within a shared visual semantic cluster be spatially adjacent in the brain, yet during visualization we consistently find contiguous patches both within and across participants. We can localize these patches of voxels to specific regions of interest (ROIs) both functionally and anatomically (See Section 6.3 for further details).
@@ 4.1 Faces
One of the first reported functionally localized areas for higher-order vision was the fusiform face area (FFA)  (Kanwisher et al., 1997). Our method identifies a face-related concept (Figure 4.1 $\varepsilon=0.55$, cluster 7) that is localized to FFA and includes voxels from all 8 participants. This cluster also has some voxels in the extrastriate body area (EBA), likely because many face images include part or all of a person’s body. We note that the positive representative images are not exclusively human and include a range of animal faces. The images often depict people eating and handling food, or holding other objects such as toothbrushes or cell phones.
@@ 4.3 Bodies
Our method identifies five notable body-related areas at $\varepsilon=0.55$ in and around EBA (downing_2001). The positive representative images for cluster 2 (Figure 5) show people and animals outside with an emphasis on legs and active movement. The negative images typically depict people indoors sitting with their legs obscured by tables. Cluster 11 (Figure 6) has a similar emphasis on hands instead of legs. People are displayed in a variety of contexts with their hands clearly visible, while the negative images are exclusively non-primate animals without hands. This cluster also shows indoor/outdoor contrast in the representative image groups and so there is some PPA activation. In Section 6.7 we outline further body-related shared voxel clusters associated with motion and a distinction between single individuals and groups of people in a sub-region of EBA.
@@ 5 Conclusion
We proposed a novel mechanism to identify shared decodable concept clusters in a large multi-participant fMRI dataset of natural image viewing (Allen et al., 2022). To achieve this, we applied a contrastive learning approach using multimodal embeddings from CLIP (Radford et al., 2021) and a novel adaptation of the DBSCAN clustering algorithm (Ester et al., 1996). Our method identifies previously reported functional areas that code for various visuo-semantic categories, as well as identifying novel shared categories. Furthermore, our method exposes interpretability by identifying the most / least associated images with CLIP-based cluster centroids. In the case of faces, we find a set of positive / negative images where the negative images depict scenarios where a face might be expected, but is not visible. This suggests that the brain responses for this SDC cluster are organized in a way that specifically codes for the presence of faces in expected visual depictions. Food is a concept that is highly confounded with other visual criteria such as shape and color. By providing the most negatively associated images for a SDC cluster, we enabled easy interpretation of semantic coding, e.g. in the case where the negated images to this SDC cluster are all grayscale images (see Section 4.2). We further identify SDC clusters specific to bodies in motion, in groups, as well as hands and legs. We identify SDC clusters related to repeated elements (numerosity), sensitivity to indoor / outdoor scenes, cardinal orientation of objects in images (horizontal / vertical) and gradations of shape information (soft vs hard) and visual perspective (close vs far). Our method showcases how various techniques from machine learning can be employed to better characterize visual representations of semantic selectivity in the human brain.
5.1 Broader Impacts
Methods that decode neural activity could have substantial impacts on both neuroscience and broader society. Identifying new category-selective brain regions could assist in diagnosing neurological and psychiatric conditions, surgical planning, development of brain-computer interfaces, and potentially helping individuals with locked-in syndrome or related disabilities to better communicate.
However, there are also significant ethical considerations with technologies that decode neural responses. If these techniques are used outside controlled research environments, there is a risk that they could be used for intrusive monitoring of mental states. As with all emerging technologies, responsible deployment is needed in order for society as a whole to maximize benefit while mitigating risks.
5.2 Limitations and Future Work
While our approach offers many advantages, it is not without drawbacks. Firstly, we are limited by the stimulus images that were chosen for the NSD experiment. Although the use of a contrastive loss function helps mitigate this, our model may still be biased toward over-represented categories in the stimulus set. Additionally, the use of CLIP may introduce its own biases, as some signal in the brain responses may not be fully captured by the CLIP embeddings. The use of fMRI also presents certain limitations. Despite its high spatial resolution, its limited temporal resolution restricts the ability to capture rapid neural activity that may play a crucial role in visual processing. Future work could address these limitations by utilizing larger and more diverse fMRI datasets, incorporating higher temporal imaging techniques such as electroencephalogram (EEG) or electrocorticography (ECoG), and exploring alternative stimulus representations beyond CLIP.
Additionally, our clustering approach might merge concepts within regions of CLIP space that have relatively uniform densities, potentially missing interesting SDCs. This issue could be addressed by adapting the heirarichcal version of DBSCAN, or by using different criteria to merge the cross-participant core points into clusters.
Furthermore, we acknowledge that our method is not a replacement for traditional hypothesis-driven experiments. It will be important for future work to test for alternative hypotheses for what might be driving responses in SDC clusters. Another interesting research direction is the application of this method to other sensory modalities beyond vision, such as auditory or somatosensory stimuli.